OpenStudy (anonymous):

anyone know what to do? i keep getting a different answer. Find the area of a sector with a central angle of 185° and a diameter of 6.4 m. Round to the nearest tenth. (1 point) 66.1 m² 5.3 m² 16.5 m² 21.0 m²

OpenStudy (mathstudent55):

Area of sector: \[A = \frac{ n ^{o} }{ 360^{o} } \pi r ^{2}\] where A = area n = measure of central angle in degrees r = radius of circle Your diameter is 6.4 m, so the radius is 3.2 m.

OpenStudy (anonymous):

10.24*(pi)*(37/72)

OpenStudy (anonymous):

is that all multiplied together to get the answer?

OpenStudy (mathstudent55):

A = (185/360) x 3.14159 x (3.2)(3.2)

OpenStudy (mathstudent55):

Multiply together: 185 x 3.14159 x 3.2 x 3.2 then divide by 360

OpenStudy (anonymous):

i got 5.1661, should i round up to 5.3...?

OpenStudy (anonymous):

answer is 16.5

OpenStudy (anonymous):

i can't get that answer no matter how i type it in to my calculator :(

OpenStudy (mathstudent55):

Type what I wrote above: 185 x 3.14 x 3.2 x 3.2 / 360

OpenStudy (anonymous):

i finally got it! thanks so much! my calculators online so i probably messed up while typing it in each time cause its pretty sensitive sorry i took so much of your time

OpenStudy (mathstudent55):

That's fine